# On the Learn-to-Optimize Capabilities of Transformers in In-Context Sparse Recovery

### v1

@@ 1 Introduction
In this paper, we examine this hypothesis through the lens of in-context sparse recovery. Sparse recovery is a classical signal processing problem that is of significant practical interest across various domains, such as compressive sensing in medical imaging (Shen et al.,, 2017) and spectrum sensing (Elad,, 2010). Compared with in-context linear regression that is often studied in the literature (Akyürek et al.,, 2022; Von Oswald et al., 2023a, ; Bai et al.,, 2024), the additional sparsity constraint makes the problem much more complicated and requires more sophisticated treatment. Our main contributions are as follows.
  • First, we demonstrate that Transformers can implement an L2O algorithm for in-context sparse recovery, and theoretically prove that a $K$-layer Transformer as an L2O algorithm can recover the underlying sparse vector in-context at a convergence rate linear in $K$. Such linear convergence results significantly improve the state-of-the-art convergence results for in-context sparse recovery (Bai et al.,, 2024; Chen et al., 2024b, ), and enrich the literature.
  • Second, we show that the Transformer as an L2O algorithm can actually outperform traditional L2O algorithms for sparse recovery in several aspects: 1) It does not require the measurement matrices involved in training to be the same as those in traditional L2O algorithms (Gregor and LeCun,, 2010; Chen et al.,, 2018; Liu and Chen,, 2019) for sparse recovery, which allows more flexibility to solve various in-context sparse recovery tasks. 2) It allows different numbers of measurements (i.e., prompt length) used for in-context sparse recovery, with guaranteed recovery performance as long as the number of measurements is sufficiently large. 3) It can extract structural properties of the underlying sparse vectors from the training data and utilize them to expedite its ICL convergence.
@@ 2 Related Works
Training Dynamics of Transformers.
There exist some works aiming to theoretically understand the ICL mechanism in Transformer through their training dynamics. Ahn et al., (2024); Mahankali et al., (2023); Zhang et al., (2023); Huang et al., (2023) investigate the dynamics of Transformers with a single attention layer and a single head for in-context linear regression tasks. Cui et al., (2024) prove that Transformers with multi-head attention layers outperform those with single-head attention. Cheng et al., (2023) show that local optimal solutions in Transformers can perform gradient descent in-context for non-linear functions. Kim and Suzuki, (2024) study the non-convex mean-field dynamics of Transformers, and Nichani et al., (2024) characterize the convergence rate for the training loss in learning a causal graph. Additionally, Chen et al., 2024a investigate the gradient flow in training multi-head single-layer Transformers for multi-task linear regression. Chen and Li, (2024) propose a supervised training algorithm for multi-head Transformers. The training dynamics of Transformers for binary classification (Tarzanagh et al., 2023b, ; Tarzanagh et al., 2023a, ; Vasudeva et al.,, 2024; Li et al.,, 2023; Deora et al.,, 2023; Li et al., 2024a, ) and next-token prediction (NTP) (Tian et al., 2023a, ; Tian et al., 2023b, ; Li et al., 2024b, ; Huang et al.,, 2024) have also been studied recently.
@@
There is a rich literature on L2O. We discuss the L2O algorithms relevant to sparse recovery here and leave more discussions on general L2O in Appendix A. Sparse recovery, typically formulated as a least absolute shrinkage and selection operator (LASSO), has many important applications like magnetic resonance imaging (Meng et al.,, 2023) and stock market forecasting (Roy et al.,, 2015), thus motivates the design of efficient algorithms. E.g., the iterative soft-thresholding algorithm (ISTA) (Daubechies et al.,, 2004) is proposed to solve LASSO and improves over the standard gradient descent algorithm. Motivated by the ISTA structure, Gregor and LeCun, (2010) introduce the Learned ISTA (LISTA), a feedforward neural network that incorporates trainable matrices into ISTA updates. Chen et al., (2018) and Liu and Chen, (2019) further propose LISTA-Partial Weight Coupling (LISTA-CP) and Analytic LISTA (ALISTA) with fewer trainable parameters, making them easier to train. They also provide theoretical analyses demonstrating a linear convergence rate. However, these existing LISTA-type algorithms only apply to a fixed measurement matrix and fail when the measurement matrices (the inputs of the input-output pairs in ICL) change. In this work, we provide evidence that Transformer can perform a LISTA-type algorithm that tackles this issue and succeeds in in-context sparse recovery tasks, both theoretically in Section 5 and experimentally in Section 6.
@@ 3.1 Transformer Architecture
where $\mathrm{mask}(\cdot)$ is defined as
$\mathrm{mask}(\mathbf{M})=\begin{bmatrix}1&1/2&\cdots&\frac{1}{N+1}\\ 0&1/2&\cdots&\frac{1}{N+1}\\ \cdots&\cdots&\cdots&\cdots\\ 0&0&\cdots&\frac{1}{N+1}\\ \end{bmatrix}\odot\mathbf{M}\quad\mbox{for}\quad\mathbf{M}\in\mathbb{R}^{(N+1)\times(N+1)},$
the operator $\odot$ denotes the Hadamard product (element-wise product), and $\sigma(\cdot)$ denotes the activation function. Let $\mathbf{h}_{i}$ denotes the $i$-th column of $\mathbf{H}$. Then, we have
$[\mathrm{Attn}_{\{(\mathbf{V}_{m},\mathbf{Q}_{m},\mathbf{K}_{m})\}}(\mathbf{H})]_{i}=\mathbf{h}_{i}+\sum_{m=1}^{M}\sum_{j=1}^{i}\frac{1}{i}\sigma((\mathbf{Q}_{m}\mathbf{h}_{i})^{\top}(\mathbf{K}_{m}\mathbf{h}_{j}))\cdot\mathbf{V}_{m}\mathbf{h}_{j}.$
@@ 4.1 In-context Sparse Recovery
In this work, we study how Transformers solve the sparse recovery problem in context. Specifically, we assume a Transformer has been pre-trained with randomly sampled sparse recovery instances. Then, during ICL, for a given sparse recovery instance with $\bm{\beta}\sim P_{\beta}$, $\{\mathbf{x}_{i}\}_{i=1}^{N}\sim P_{\mathbf{x}}$, $\{\epsilon_{i}\}_{i=1}^{N}\sim P_{\epsilon}$, the Transformer aims to predict $y_{N+1}$ with input $(\mathbf{X},\mathbf{y},\mathbf{x}_{N+1})$, where $y_{N+1}=\mathbf{x}_{N+1}^{\top}\bm{\beta}+\epsilon_{N+1}$.
@@ 4.3 Transformer Can Provably Perform LISTA-type Algorithms
where $\{\bm{\beta}_{i}^{(1)}\}_{i\in[2N+1]}\in\mathbb{R}^{d}$ are implicit parameter vectors initialized as $\mathbf{0}_{d}$, and $\mathbf{x}_{i}$ is the $i$-th column of the transposed measurement matrix, i.e, $[\mathbf{X}^{\top}]_{:,i}$.
@@ 5.1 Sparse Vector Estimation
Let $\delta\in(0,1)$, $N_{0}=8(4S-2)^{2}{\frac{\log d+\log S-\log\delta}{c}}$, $\alpha_{n}=-\log\big{(}1-\frac{2}{3}\gamma+\gamma(2S-1)\sqrt{\frac{\log d-\log\delta}{nc}}+\sqrt{\frac{\log S-\log\delta}{nc}}\big{)}$, where $c$ is a positive constant and $\gamma$ is a positive constant satisfies $\gamma\leq\frac{3}{2}$. For a $K$-layer Transformer model with the structure described in Section 4.3, under Assumption 1, there exists a set of parameters such that for any $n\in[N_{0}:N]$, with probability at least $1-\delta$, we have
@@ 5.2 Label Prediction
Under the same setting as in Section 4.3, with probability at least $1-n\delta-\delta^{\prime}$, we have
@@
for a linear read-out function, and with probability at least $1-\delta$, we have
$\|y_{n}-\widehat{y}_{n}\|\leq c_{5}e^{-\alpha_{n}K}$
for a query read-out function, where $c_{4}$, $c_{5}$ are constants.
@@ 6 Experimental Results
Baselines. The baselines for our experiments include traditional iterative algorithms such as ISTA and FISTA (Beck and Teboulle,, 2009). We also evaluate three classical LISTA-type L2O algorithms: LISTA, LISTA-CP, ALISTA (Gregor and LeCun,, 2010; Chen et al.,, 2018; Liu and Chen,, 2019). For all of these algorithms, we set the number of iterations $K=12$. We generate a single fixed measurement matrix $\mathbf{X}$. For each training epoch, we create $N=50,000$ instances from 50,000 randomly generated sparse vectors. We also evaluate the LISTA-VM algorithm introduced in Theorem 4.1, where we set the number of iterations $K=12$ as well. For each training epoch, we randomly sample 100 measurement matrices, each generating 500 instances from 500 randomly generated sparse vectors, which results in a total of 50,000 instances. For comparison, we also meta-train LISTA and LISTA-CP using the same training method as LISTA-VM. We do not perform meta-training for ALISTA, as the training process of ALISTA involves solving a non-convex optimization problem for each different measurement matrix $\mathbf{X}$, which makes meta-training for ALISTA unrealistic. For all baseline algorithms, we minimize the sparse vector prediction loss $\sum_{j=1}^{N}\|\widehat{\bm{\beta}}_{j}-\bm{\beta}_{j}\|^{2}$ using gradient descent for each epoch. We run all baseline experiments for 340 epochs.
Transformer structure. We consider two Transformer models, i.e., a small Transformer model (denoted as Small TF) and GPT-2. Small TF has $12$ layers, each containing a self-attention layer followed by an MLP layer. Each self-attention layer has 4 attention heads. We set the embedding dimension to $D=42$, and the embedding structure according to Equation 4.2. For GPT-2, we employ $12$ layers, and set the embedding dimension to $256$ and the number of attention heads per layer to $8$. In order to train Small TF and GPT-2, we randomly generate $64$ instances per epoch and train the algorithms for $10^{6}$ epochs. The training process minimizes the label prediction loss $\sum_{j=1}^{T}(y_{N+1,j}-\widehat{y}_{N+1,j})^{2}$. We run the experiments for Small TF and GPT-2 on an NVIDIA RTX A5000 GPU with 24G memory. The training time for Small TF is approximately $8$ hours, while the training time for GPT-2 is around $12$ hours.
@@
We first do not impose any support constraint on $\bm{\beta}$. We start with the general setting where the testing instance is randomly generated (Varying $\mathbf{X}$). As shown in Figure 1(a), GPT-2 outperforms Small TF, followed by LISTA-VM, which outperforms iterative algorithms FISTA and ISTA, while the classical LISTA-type algorithms LISTA, LISTA-CP, and ALISTA perform the worst. Such results highlight the efficiency of LISTA-VM, Small TF and GPT-2 in solving ICL sparse recovery problems, corroborating our theoretical result in Theorem 5.1. Meanwhile, classical LISTA-type algorithms cannot handle mismatches between the measurement matrices in pre-training and testing, leading to poor prediction performance. When $\mathbf{X}$ during testing is fixed to be the same as that in pre-training, all LISTA-type algorithms achieve performances comparable with Small TF. For meta-trained LISTA and LISTA-CP, the corresponding prediction loss (denoted by red marks in Figure 1(a)) are still much higher than that under LISTA-VM. This indicates that meta-training cannot help those classical LISTA-type algorithms achieve performances comparable with LISTA-VM, further corroborating the strong generalization capability induced by the constructed Transformer structure.
Next, we impose an additional constraint on the support of the sparse vectors. For in-context sparse recovery instances used in both pre-training and testing, we set the support of the sparse vector $\bm{\beta}$ to the first $10$ entries, i.e., $\mathbb{S}=\{1,2,\cdots,10\}$. As observed in Figure 1(b), Small TF and GPT-2 significantly improve their performances in Figure 1(a), while other baseline algorithms do not exhibit significant performance improvements. In Figure 1(b), we present the experimental results for a support-selected version of LISTA-VM, referred to as LISTA-VM-SS. This algorithm is a simple variation of LISTA-VM, where we incorporate prior knowledge of the support by setting all columns in $\mathbf{X}$ whose indices are not in the prior support set to be zero vectors. As we claim in Remark 3 and Corollary D.1, a Transformer could perform this LISTA-VM-SS by utilizing prior knowledge of the support. Our results show that the LISTA-VM-SS achieves an in-context prediction error of approximately $0.07$, which is almost 10 times better than the standard version of LISTA-VM and is comparable to the prediction error of Small TF. This empirical finding corroborates our claim in Remark 3.
@@ 7 Conclusion
In this work, we demonstrated that Transformers’ known ICL capabilities could be understood as performing L2O algorithms. Specifically, we showed that for sparse recovery tasks, Transformers can execute the LISTA-VM algorithm with a provable linear convergence rate. Our results highlight that, unlike existing LISTA-type algorithms, which are limited to solving individual sparse recovery problems with fixed measurement matrices, Transformers can address a general class of sparse recovery problems with varying measurement matrices without requiring parameter updates. Experimentally, we demonstrated that Transformers can leverage prior knowledge from training tasks and generalize effectively across different lengths of demonstration pairs, where traditional L2O methods typically fail. These insights not only deepen our understanding of Transformers’ capabilities in ICL but also suggest new potential applications for utilizing Transformers in other optimization tasks.

### v2

@@ 1 Introduction
In this paper, we examine this hypothesis through the lens of in-context sparse recovery. Sparse recovery is a classical signal processing problem that is of significant practical interest across various domains, such as compressive sensing in medical imaging (Shen et al.,, 2017) and spectrum sensing (Elad,, 2010). Recent works show that Transformers are able to implement gradient descent-based algorithms with sublinear convergence rates for in-context sparse recovery (Bai et al.,, 2024; Chen et al., 2024b, ). However, empirical findings indicate that Transformers can solve in-context sparse recovery more efficiently than gradient descent-based approaches (Bai et al.,, 2024). Meanwhile, there exists a plethora of L2O algorithms that solve the classical sparse recovery problem efficiently with linear convergence guarantees (Gregor and LeCun,, 2010; Chen et al.,, 2018; Liu and Chen,, 2019). Therefore, examining the L2O capabilities of Transformers in solving the in-context sparse recovery task becomes a promising direction and may serve as a perfect example to validate our hypothesis. Our main contributions are as follows.
  • First, we demonstrate that Transformers can implement an L2O algorithm for in-context sparse recovery, and theoretically prove that a $K$-layer Transformer as an L2O algorithm can recover the underlying sparse vector in-context at a convergence rate linear in $K$. The linear convergence results in this work significantly improve the state-of-the-art convergence results for in-context sparse recovery and validate our previous hypothesis.
  • Second, we show that the Transformer as an L2O algorithm can actually outperform traditional L2O algorithms for sparse recovery in several aspects: 1) It does not require the measurement matrices involved in training to be the same, which is in stark contrast to traditional L2O algorithms (Gregor and LeCun,, 2010; Chen et al.,, 2018; Liu and Chen,, 2019) for sparse recovery, and allows more flexibility to solve various in-context sparse recovery tasks. 2) It allows different numbers of measurements (i.e., prompt length) used for in-context sparse recovery, with guaranteed recovery performance as long as the number of measurements is sufficiently large. 3) It can extract structural properties of the underlying sparse vectors from the training data and utilize them to expedite its ICL convergence.
@@ 2 Related Works
The L2O framework, as summarized in the review paper (Chen et al.,, 2022), is an optimization paradigm that develops an optimization method (i.e., a solver) by training across a set of similar problems (tasks) sampled from a task distribution. While the training process is often offline and time-consuming, the objective of L2O is to improve the optimization efficiency and accuracy when the method is deployed online and any new task sampled from the same distribution is encountered.
As a general optimization framework, L2O has demonstrated its advantage over classical static optimization frameworks in various optimization problems and applications. In this work, we focus on the L2O algorithms relevant to sparse recovery here and leave more discussions on general L2O in Appendix A. Sparse recovery, typically formulated as a least absolute shrinkage and selection operator (LASSO), has many important applications like magnetic resonance imaging (Meng et al.,, 2023) and stock market forecasting (Roy et al.,, 2015), thus motivates the design of efficient algorithms. E.g., the iterative soft-thresholding algorithm (ISTA) (Daubechies et al.,, 2004) is proposed to solve LASSO and improves over the standard gradient descent algorithm. Motivated by the ISTA structure, Gregor and LeCun, (2010) introduce the Learned ISTA (LISTA), a feedforward neural network that incorporates trainable matrices into ISTA updates. Chen et al., (2018) and Liu and Chen, (2019) further propose LISTA-Partial Weight Coupling (LISTA-CP) and Analytic LISTA (ALISTA) with fewer trainable parameters, making them easier to train. They also provide theoretical analyses demonstrating a linear convergence rate.
We note that for these existing LISTA-type algorithms, both the training and testing tasks (instances) are randomly generated by fixing the measurement matrix but varying the underlying sparse vectors. In general, they cannot handle cases where the training instances are generated with varying measurement matrices, or the measurement matrices during training and testing do not match.
@@ 3.1 Transformer Architecture
where $\mathrm{mask}(\mathbf{M})$ satisfies $\left[\mathrm{mask}(\mathbf{M})\right]_{i,j}=\frac{1}{j}\mathbf{M}_{i,j}$ if $i\leq j$ and $\left[\mathrm{mask}(\mathbf{M})\right]_{i,j}=0$ otherwise. $\sigma(\cdot)$ denotes the activation function.
@@ 4.1 In-context Sparse Recovery
In this work, we study how Transformers solve the sparse recovery problem in context. For the pre-training process, a set of in-context sparse recovery instances $\{(\mathbf{X}^{j},\mathbf{y}^{j},\mathbf{x}_{N+1}^{j},y_{N+1}^{j})\}_{j=1}^{N_{\text{train}}}$ is generated according to the relationship $y_{n}^{j}=(\mathbf{x}_{n}^{j})^{\top}\bm{\beta}^{j}+\epsilon_{n}^{j}$ for $j\in[N_{\text{train}}]$ and $n\in[N+1]$, where $\bm{\beta}^{j}\sim P_{\beta}$, $\mathbf{x}_{n}^{j}\sim P_{\mathbf{x}}$, and $\epsilon_{n}^{j}\sim P_{\epsilon}$, respectively. Given the set of training instances, a pre-trained Transformer is obtained by minimizing a certain loss function.
After pre-training, during the inference process for ICL, a sparse recovery instance $(\mathbf{X},\mathbf{y},\mathbf{x}_{N+1})$ is sampled randomly according to the same distributions as in the pre-training, and the Transformer then aims to predict $y_{N+1}$ using the input $(\mathbf{X},\mathbf{y},\mathbf{x}_{N+1})$ without any further parameter updating.
@@ 4.3 Transformer Can Provably Perform LISTA-type Algorithms
where $\{\bm{\beta}_{i}^{(1)}\}_{i\in[2N+1]}\in\mathbb{R}^{d}$ are implicit parameter vectors initialized as $\mathbf{0}_{d}$, and $\mathbf{x}_{i}$ is the $i$-th column of the transposed measurement matrix, i.e, $[\mathbf{X}^{\top}]_{:,i}$. We note that a similar embedding structure is adopted in Bai et al., (2024).
@@ 5.1 Sparse Vector Estimation
Let $\delta\in(0,1)$, $N_{0}=8(4S-2)^{2}{\frac{\log d+\log S-\log\delta}{c}}$, $\alpha_{n}=-\log\big{(}1-\frac{2}{3}\gamma+\gamma(2S-1)\sqrt{\frac{\log d-\log\delta}{nc}}+\sqrt{\frac{\log S-\log\delta}{nc}}\big{)}$, where $c$ is a positive constant and $\gamma$ is a positive constant satisfies $\gamma\leq\frac{3}{2}$. For a $K$-layer Transformer model with the structure described in Section 4.3, under Assumption 1, there exists a set of parameters such that for any randomly generated sparse recovery instance and any $n\in[N_{0}:N]$, with probability at least $1-\delta$, we have
@@
Remark 2 (Linear convergence rate).
The linear convergence rate demonstrated in Theorem 5.1 is due to the incorporation of curvature information into the update rule. Specifically, the learned matrices $\mathbf{M}^{(k)}$ serve as approximations of the inverse Hessian. By leveraging the statistical properties of the problem, these matrices effectively accelerate convergence, enabling the Transformer to mimic second-order optimization methods. This allows the Transformer to overcome the traditional limitations of first-order methods, which are typically restricted to a sublinear $\mathcal{O}(1/K)$ convergence rate unless additional assumptions are introduced (Nesterov,, 2005; Nemirovskij and Yudin,, 1983).
@@ 5.2 Label Prediction
Under the same setting as in Section 4.3, for any $n\in[N_{0}+1:N+1]$, with probability at least $1-n\delta-\delta^{\prime}$, we have
@@
for a linear read-out function, and with probability at least $1-\delta$, we have $\|y_{n}-\widehat{y}_{n}\|\leq c_{5}e^{-\alpha_{n}K}$ for a query read-out function, where $c_{4}$, $c_{5}$ are constants.
@@ 6 Experimental Results
Baselines. The baselines for our experiments include traditional iterative algorithms such as ISTA and FISTA (Beck and Teboulle,, 2009). We also evaluate three classical LISTA-type L2O algorithms: LISTA, LISTA-CP, ALISTA (Gregor and LeCun,, 2010; Chen et al.,, 2018; Liu and Chen,, 2019). For all of these algorithms, we set the number of iterations $K=12$. We generate a single fixed measurement matrix $\mathbf{X}$. For each training epoch, we create $I=50,000$ instances from $50,000$ randomly generated sparse vectors. During inference, we evaluate the pre-trained LISTA-type models under two settings: (1) when the measurement matrix remains identical to that used during pre-training, reported as “Fixed $X$”, and (2) when the measurement matrix is varying through random sampling, reported as “Varying $X$”.
We also evaluate the LISTA-VM algorithm introduced in Theorem 4.1, where we set the number of iterations $K=12$ as well. For each training epoch, we randomly sample $100$ measurement matrices, each generating $500$ instances from 500 randomly generated sparse vectors, which results in a total of $50,000$ instances. For comparison, we also meta-train LISTA and LISTA-CP using the same training method as LISTA-VM. We do not perform meta-training for ALISTA, as the training process of ALISTA involves solving a non-convex optimization problem for each different measurement matrix $\mathbf{X}$, which makes meta-training for ALISTA unrealistic. For all baseline algorithms, we minimize the sparse vector prediction loss $\sum_{i=1}^{I}\|\widehat{\bm{\beta}}_{i}-\bm{\beta}_{i}\|^{2}$ using gradient descent for each epoch. We run all baseline experiments for 340 epochs.
Transformer structure. We consider two Transformer models, i.e., a small Transformer model (denoted as Small TF) and GPT-2. Small TF has $12$ layers, each containing a self-attention layer followed by an MLP layer. Each self-attention layer has 4 attention heads. We set the embedding dimension to $D=42$, and the embedding structure according to Equation 4.2. For GPT-2, we employ $12$ layers, and set the embedding dimension to $256$ and the number of attention heads per layer to $8$. We note that the Small TF model shares the same 4-head configuration and ReLU activation function as described in Theorem 4.1, while the configuration of GPT-2 is commonly used in practical applications. In order to train Small TF and GPT-2, we randomly generate $64$ instances per epoch and train the algorithms for $10^{6}$ epochs. The training process minimizes the label prediction loss $\sum_{j=1}^{N+1}(y_{j}-\widehat{y}_{j})^{2}$. We run the experiments for Small TF and GPT-2 on an NVIDIA RTX A5000 GPU with 24G memory. The training time for Small TF is approximately $8$ hours, while the training time for GPT-2 is around $12$ hours.
@@
We first do not impose any support constraint on $\bm{\beta}$. We start with the general setting where the testing instance is randomly generated (Varying $\mathbf{X}$). As shown in Figure 1(a), GPT-2 outperforms Small TF, followed by LISTA-VM, which outperforms iterative algorithms FISTA and ISTA, while the classical LISTA-type algorithms LISTA, LISTA-CP, and ALISTA perform the worst. Such results highlight the efficiency of LISTA-VM, Small TF and GPT-2 in solving ICL sparse recovery problems, corroborating our theoretical result in Theorem 5.1. Meanwhile, classical LISTA-type algorithms cannot handle mismatches between the measurement matrices in pre-training and testing, leading to poor prediction performance. When $\mathbf{X}$ during testing is fixed to be the same as that in pre-training, all LISTA-type algorithms achieve performances comparable with Small TF. For meta-trained LISTA and LISTA-CP, the corresponding prediction loss (denoted by red marks in Figure 1(a)) are still much higher than that under LISTA-VM. This indicates that meta-training cannot help those classical LISTA-type algorithms achieve performances comparable with LISTA-VM, further corroborating the strong generalization capability induced by the constructed Transformer.
Next, we impose an additional constraint on the support of the sparse vectors. For in-context sparse recovery instances used in both pre-training and testing, we set the support of the sparse vector $\bm{\beta}$ to the first $10$ entries, i.e., $\mathbb{S}=\{1,2,\cdots,10\}$. As observed in Figure 1(b), Small TF and GPT-2 significantly improve their performances in Figure 1(a), while other baseline algorithms do not exhibit significant performance improvements. In Figure 1(b), we present the experimental results for a support-selected version of LISTA-VM, referred to as LISTA-VM-SS. This algorithm is a simple variation of LISTA-VM, where we incorporate prior knowledge of the support by setting all columns in $\mathbf{X}$ whose indices are not in the prior support set to be zero vectors. As we claim in Remark 4 and Corollary D.1, a Transformer could perform this LISTA-VM-SS by utilizing prior knowledge of the support. Our results show that the LISTA-VM-SS achieves an in-context prediction error of approximately $0.07$, which is almost 10×10$\times$10 × better than the standard version of LISTA-VM and is comparable to the prediction error of Small TF. This empirical finding corroborates Remark 4.
@@ 7 Conclusion
In this work, we demonstrated that Transformers’ known ICL capabilities could be understood as performing L2O algorithms. Specifically, we showed that for in-context sparse recovery tasks, Transformers can execute the LISTA-VM algorithm with a provable linear convergence rate. Our results highlight that, unlike existing LISTA-type algorithms, which are limited to solving individual sparse recovery problems with fixed measurement matrices, Transformers can address a general class of sparse recovery problems with varying measurement matrices during inference without requiring parameter updates. Experimentally, we demonstrated that Transformers can leverage prior knowledge from training tasks and generalize effectively across different lengths of demonstration pairs, where traditional L2O methods typically fail.
Acknowledgments
The work of R. Liu and J. Yang was supported in part by the U.S. National Science Foundation under the grants ECCS-2133170 and ECCS-2318759. The work of R. Zhou was supported in part by NSF grants 2139304, 2146838 and the Army Research Laboratory grant under Cooperative Agreement W911NF-17-2-0196. The work of C. Shen was supported in part by the U.S. National Science Foundation under the grants CNS-2002902, ECCS-2029978, ECCS-2143559, ECCS-2033671, CPS-2313110, and ECCS-2332060.